OpenStudy (anonymous):

Given the parent function of f(x) = x4, what change will occur when the function is changed to -f(1/2x) ? Graph opens the same way and is narrower Graph opens the same way and is wider Graph opens the opposite way and is narrower Graph opens the opposite way and is wider

mathslover (mathslover):

Is it f(x) = x^4 or f(x) = x4 \(x^4 \ \text{OR}\ x4 \ ? \)

OpenStudy (anonymous):

sorry yea its x^4

mathslover (mathslover):

Fine. So, you need to find \[ - f\left( \cfrac{1}{2x} \right) \] Right?

OpenStudy (anonymous):

thats right just the x is right next to the line not on the bottom with the 2

mathslover (mathslover):

Oh Okay. That will make a bit difference in the final answer. Now in order to find the above function in (1/ (2) * x ) .. then just plugin \( \cfrac{1}{2} x \) in place of \(x\) in the parent function i.e. \(f(x) = x^4 \)

mathslover (mathslover):

For example, if you are require to find f(2) then you will just plugin x = 2 in the parent function i.e. \[ \boxed{ \mathsf{ f(x) = x^4 \\\implies f(2) = 2^4 \\ \implies f(2) = 16 } } \] So, this is the way you will have to find \( - f\left( \cfrac{1}{2} x \right) \) Got it?

OpenStudy (anonymous):

yeah kinda so \[1/2^{4}\] ?

mathslover (mathslover):

Not actually as you've ignored "x" there.

mathslover (mathslover):

See you have to plug in \(\cfrac{1}{2} \times x \) at the place of just \(x\) in the parent function. So, you will get something like thus: \({ f\left( \cfrac{1}{2} x \right) = \left( \cfrac{1}{2} x \right)^4 } \)

mathslover (mathslover):

Now, if you will solve it further, you will get: \(\left( \cfrac{1 \times x^4 }{ 2^4} \right) \) which is equivalent to \(\left( \cfrac{x^4}{2^4} \right) \) Right?

mathslover (mathslover):

See, why am I doing this all? Because this will help us to notice the changes in the graphs. I need you to give me some faster responses as I'm not here for much time.

OpenStudy (anonymous):

oh ok verry sorry i had to do something. but yes i am follwing what your saying

mathslover (mathslover):

Okay, now, notice that what I have found is \(f \left( \cfrac{1}{2} x \right)\) , that is I have not taken negative till now. Now, just take the negative of \(\cfrac{x^4}{16}\) that is you get: \(- f \left( \cfrac{1}{2} x \right) = - \cfrac{x^4}{16} \) \(\color{blue}{\text{Originally Posted by}}\) @mathslover See you have to plug in \(\cfrac{1}{2} \times x \) at the place of just \(x\) in the parent function. So, you will get something like thus: \({\color{blue}{ f\left( \cfrac{1}{2} x \right)} = \left( \cfrac{1}{2} x \right)^4 } \) \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

ohh ok i get the steps now. so the graph opens the same way and wider

mathslover (mathslover):

Yeah, right. http://www.wolframalpha.com/input/?i=x%5E4+and+-+x%5E4+%2F+16 Notice the two graphs there, the graph of -x^4/16 will open the opposite way and wider (because of that 16 in the denominator)

mathslover (mathslover):

I hope this helped. Good Luck!

OpenStudy (anonymous):

yes you help verry much. perfect job at explaining, thank you! (:

mathslover (mathslover):

Good to hear that Andy. Take care of yourself and good luck for future studies. I wish you lots of good luck for your journey as a student all over the life as well as the journey on OpenStudy! (:

OpenStudy (anonymous):

thank you that was verry nice. i wish you the same! (:

OpenStudy (anonymous):

This is wrong, at least if you meant f(x)= x^4. (just took the test)